OpenStudy (anonymous):

@mlamott15 Math Help!!!

OpenStudy (anonymous):

haha i feel so special XD whats up tho??

OpenStudy (anonymous):

Find the inner product of vector a and vector b if vector a = (4, -2, -2) and vector b = (-7, -2, 4), and state whether the vectors are perpendicular.

OpenStudy (anonymous):

dude like i said....i had someone do my vector homework for me so i have no clue

OpenStudy (anonymous):

:( haha what about @chaser71 or @ilovehim14<3

OpenStudy (anonymous):

dude @chaser71 dosent know :P why ask him?? hahaha

OpenStudy (anonymous):

hey dont forget me ask @Twilkis12 lol he might

OpenStudy (anonymous):

Hey @Twilkis12

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

ugh! no!!! lol

OpenStudy (anonymous):

Find the inner product of vector a and vector b if vector a = (4, -2, -2) and vector b = (-7, -2, 4), and state whether the vectors are perpendicular.

OpenStudy (anonymous):

@Destinymasha @RadEn @eliassaab @phi @wio @ParthKohli @campbell_st @AravindG @Luigi0210 @2epdrix @tkhunny

OpenStudy (tkhunny):

Well, do it. 4*(-7) + (-2)*(-2) + (-2)*4 = ?? And, what does that tell us?

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

-32 @tkhunny

OpenStudy (anonymous):

What does that mean?

OpenStudy (tkhunny):

Okay, now the second question. What does that tell us? Perpendicular implies an inner product of zero (0). Did we get zero (0)?

OpenStudy (anonymous):

no. So, if it is not 0 it is not perpendicular?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Why is there still a question? Go with confidence. Zero? Perpendicular. Not Zero? Not Perpendicular.

OpenStudy (anonymous):

Help with one more? A triangular prism has vertices at A(2, 0, 0), B(2, 1, 3), C(2, 2, 0), D(0, 0, 0), E(0, 1, 3), and F(0,2, 0). Which image point has the coordinates (1, 4, 3) after a translation using the vector (1, 2, 3)?

OpenStudy (anonymous):

@tkhunny